# Policy Optimization in RLHF: The Impact of Out-of-preference Data

### 17 Dec 2023 v1

@@ 3 Experiments
In this section, we conduct numerical experiments to validate the improvement of RMB-PO and RMB-PO+ by better stochastic approximation. All of our experiments are run with 10 different random seeds (2021-2030), and the averaged results are reported. Note that we set $\pi_{\text{REF}}$ to be a policy with a uniform action distribution in all experiments and $\beta=0.01$ for all methods. Besides, we use a policy with a uniform action distribution to collect the preference data.
@@ 3.1 Linear Bandit
In our experiments, we set the size of preference data to be $n=20$ and the size of preference-free data to be $m=10n$, resulting in training accuracy of the reward model ranging from 60% to 80%. We display the optimality gap $|r(\pi^{\star})-r(\widehat{\pi})|$ (the smaller, the better) in Figure 2 and Figure 2, where $r(\pi)$ is the expected performance, i.e., $r(\pi)=\mathbb{E}_{s\sim\rho(\cdot)}\mathbb{E}_{a\sim\pi(\cdot|s)}[r(s,a)]$ (in our experiments, we use 5000 sampled states to approximate this expectation).
@@ 4 Discussion
Our results are related to imitation learning (Osa et al., 2018), where the goal is to learn a policy from expert demonstrations. A popular approach to achieve this goal is through behavioral cloning (BC) (Pomerleau, 1991), which trains a policy model by maximizing the likelihood of expert data. Note that the working mechanism of BC is quite similar to DPO, as in Equation 5, where the likelihood of positively preferred actions is increased and that of negatively preferred actions is decreased:
@@
where $D_{\operatorname{E}}$ is the expert dataset. In contrast, Ghasemipour et al. (2019) showed that adversarial imitation learning (AIL) methods, such as GAIL (Ho and Ermon, 2016), leverage a recovered reward function to perform policy optimization on “out-of-expert-data” through online interaction, significantly improving performance. Following the formulation in (Xu et al., 2022), the training objective of reward-model-based AIL can be re-formulated as
@@
where $\widehat{d}_{\operatorname{E}}$ is the empirical state-action distribution estimated from $D_{\operatorname{E}}$, and $d_{\pi}(s,a)$ is obtained from online interaction. For the above optimization problem, states beyond those in the expert dataset are used. We notice that Xu et al. (2022) theoretically proved that AIL, by leveraging additional online data for imitation, can outperform BC.
@@
Finally, we note that compared with reward-model-free methods such as DPO (Rafailov et al., 2023), reward-model-based policy optimization (RMB-PO) methods do not require extra preference annotation. For applications such as large language models (LLMs), training and storing a reward model has been shown to be highly efficient, as demonstrated in (Yao et al., 2023). The primary challenge in RMB-PO lies in the huge action space during policy optimization. However, this issue can be effectively addressed by computationally efficient methods like those proposed by (Dong et al., 2023; Li et al., 2023).
@@ 5 Conclusion
Our results underscore the importance of policy optimization on out-of-preference data to unlock the reward’s generalization capacity. Otherwise, the policy may conflict with the learned reward model on out-of-preference data and suffer a poor performance.

### 25 Feb 2024 v2

@@ 3 Experiments
In this section, we conduct numerical experiments to validate the improvement of RMB-PO and RMB-PO+ by better stochastic approximation. All of our experiments are run with 10 different random seeds (2021-2030), and the averaged results are reported [^2] We exclude the worst and best results to make a robust estimation of the performance.. Note that we set $\pi_{\text{REF}}$ to be a policy with a uniform action distribution in all experiments and $\beta=0.01$ for all methods. Besides, we use a policy with a uniform action distribution to collect the preference data.
@@ 3.1 Linear Bandit
We believe that in scenarios where $\phi_{\pi}\neq\phi_{r}$, RMB-PO approaches could exhibit more promising performance than RMF-PO approaches. This is because, in such cases, the policy and reward models may align well by learning from preference data. However, in out-of-preference-distribution scenarios, they may extrapolate and generalize quite differently due to mismatches in representations. Nevertheless, RMB-PO approaches could use out-of-preference-distribution data to mitigate these mismatches and tend to perform well. The case where $\phi_{\pi}\neq\phi_{r}$ will be revisited in later neural bandit experiments, where the policy model and reward model typically utilize distinct architectures and learn distinct representations.
In our experiments, we set the size of preference data to be $n=20$ and the size of preference-free data to be $m=10n$, resulting in training accuracy of the reward model ranging from 60% to 80% over 10 experiments. We display the optimality gap $|r(\pi^{\star})-r(\widehat{\pi})|$ (the smaller, the better) in Figure 3 and Figure 3, where $r(\pi)$ is the evaluation performance of a policy $\pi$, i.e., $r(\pi)=\mathbb{E}_{s\sim\rho(\cdot)}\mathbb{E}_{a\sim\pi(\cdot|s)}[r(s,a)]$ (in our experiments, we use 5000 sampled states to approximate this expectation).
@@
To gain a better understanding, we also visualize the learned policy distribution in the $\phi_{\pi}\neq\phi_{r}$ setting; see Figure 4. To observe the training distribution coverage, we plot the states from the preference dataset. Additional states used in RMB-PO+ almost cover the entire state space but are not shown for readability reasons. From the reported curves, we observe that DPO aligns well with the optimal policy in the regions covered by preference data, and RMB-PO(+) methods tend to perform better than DPO in the out-of-distribution regime not covered by the preference data.
(a) Action $a_{0}$.
(b) Action $a_{1}$.
(c) Action $a_{2}$.
(d) Action $a_{3}$.
Figure 4: Probabilities of four actions $a_{0}$, $a_{1}$, $a_{2}$ and $a_{3}$. Results illustrate that RMB-PO(+) methods leverage out-of-preference data to better learn the policy distribution on out-of-distribution states and improve the generalization performance.
@@ 4 Discussion
Our research is related to imitation learning (Osa et al., 2018), which aims to learn a policy from expert demonstrations. A popular approach to achieve this goal is through behavioral cloning (BC) (Pomerleau, 1991), which trains a policy model by maximizing the likelihood of expert data. Note that the working mechanism of BC is quite similar to DPO, as in Equation 4, where the likelihood of positively preferred actions is increased and that of negatively preferred actions is decreased:
@@
where $D_{\operatorname{E}}$ is the expert dataset.
Ghasemipour et al. (2019) showed that another class of imitation methods, known as adversarial imitation learning (AIL) methods, (such as GAIL (Ho and Ermon, 2016)), usually performs better than BC. In particular, AIL methods leverage a recovered reward function to perform policy optimization on “out-of-expert-data” through online interaction, significantly improving performance. Following the formulation in (Xu et al., 2022), the training objective of reward-model-based AIL can be re-formulated as
@@
where $\widehat{d}_{\operatorname{E}}$ is the empirical state-action distribution estimated from $D_{\operatorname{E}}$, and $d_{\pi}(s,a)$ is obtained from online interaction. For the optimization objective of AIL, it utilizes states beyond those in the expert dataset (reflected in the summation over all state-action pairs). We notice that Xu et al. (2022) theoretically proved that AIL can outperform BC in terms of addressing the distribution shift issue with optimization on “out-of-expert data”. The idea of recovering a reward function and using it to perform extensive policy optimization is quite similar to the framework of RLHF.
@@
Finally, we note that compared with reward-model-free methods such as DPO (Rafailov et al., 2023), reward-model-based policy optimization (RMB-PO) methods do not require extra preference annotation. For applications such as language models, training and storing a reward model has been shown to be highly efficient, as demonstrated in (Yao et al., 2023). The primary challenge in RMB-PO lies in the huge action space during policy optimization. However, this issue can be effectively addressed by computationally efficient methods like those proposed by (Dong et al., 2023; Li et al., 2023). Notably, Li et al. (2023) showed that optimizing the language model with prompts-only data can improve performance, a setting that cannot achieved by reward-model-free approaches such as DPO.
@@ 5 Conclusion
We analyze the errors of policy optimization methods when learning from preferences for alignment. We also conduct experiments to validate our claims. Our results underscore the importance of optimizing policies on out-of-preference data and the power of using a reward model to provide supervision signals.